OpenStudy (deadshot):

What is the slope of the line through the points (5, 7) and (4, 7)?

hartnn (hartnn):

use the formula for slope (m) \(\large m=\dfrac{y_2-y_1}{x_2-x_1}\)

hartnn (hartnn):

you know what x1,y1 and x2,y2 are ?

OpenStudy (anonymous):

Use this formula : m=y1-y2/x1-x2 y-y1=m(x-x1)

OpenStudy (deadshot):

x1 and y1 are (5,7), and x2 and y2 are (4,7), right?

OpenStudy (anonymous):

Yeah !

hartnn (hartnn):

correct! x1 = 5 y1=7 x2=4 y2=7 just plug in values :)

hartnn (hartnn):

what do you get for m ?

OpenStudy (deadshot):

0/1

hartnn (hartnn):

which is 0, right ? correct, your m= slope = 0 :)

OpenStudy (deadshot):

Thanks!

hartnn (hartnn):

welcome ^_^